# Exploring Exponential Runge-Kutta Methods: A Survey

### v2

@@ 2.3. Stability
A good starting point is provided by the paper published in 2009 by Maset and Zennaro [maset2009unconditional], in which the authors defined the concepts of unconditional stability for ExpRK methods. Their definition is based on the concepts of contractivity and asymptotic stability for semilinear systems of the form (1) where $g$ is non-stiff. Specifically, given solutions $u_{1}$, $u_{2}$ of (1) and $\delta:=u_{1}-u_{2}$, (1) is called contractive if
@@
Maset and Zennaro [maset2009unconditional] obtained sufficient conditions for unconditional contractivity and asymptotic stability in terms of a suitable stability function. This allows them to conclude that the ETD Euler method (3) and the two main classes of ExpRK methods for certain values of the parameters are unconditionally contractive and asymptotically stable. In 2013, the same authors [maset2013stability] sharpened the aforementioned results and determined necessary and sufficient conditions for unconditional contractivity and asymptotic stability. This implies that no other ExpRK method with 2 stages, other than those already detected in [maset2009unconditional], is unconditionally stable. They also analyzed the main methods with 3 to 5 stages, and observed that none of them is unconditionally stable either. However, since $\gamma$ is assumed to be relatively small, conditional contractivity and stability properties are also desirable.
@@
Sufficient conditions for conditional properties are provided in [maset2013stability]. Without any restriction on the norm, these conditions only guarantee conditional contractivity [asymptotic stability] for those methods which are already known to be unconditionally contractive [asymptotically stable]. However, more can be said if the Euclidean norm is considered. In particular, weaker restrictions on the parameters of two-stages methods are needed in order to obtain conditional properties. The higher-order methods considered do not satisfy the sufficient conditions for general problems of the form (1), even when considering the Euclidean norm, but many satisfy conditional stability properties when restricted to certain classes of problems, such as those where AA is symmetric.
@@ 3. The historical path of ExpRK methods
In this section, we present a brief historical overview of the development and dissemination of ExpRK methods, tracing their evolution from their inception to the 2000s. We identified two major periods in which to structure our analysis. The first period begins in the late 1950s and lays the theoretical foundations that would gain traction in the following decades, up until the second half of the 1970s. The second period marks a resurgence of interest in these methods, starting in the late 1980s. Several works have already synthesized the progress made up to the early 2000s (see e.g. [minchev2005review]). Therefore, we will not dwell on the details and instead refer the reader to the relevant literature. Although the 2005 works by Hochbruck and Ostermann [hochbruck2005exponential, hochbruck2005explicit] represent a significant milestone in the evolution of ExpRK methods—marking the first application of these methods to semilinear parabolic problems—we have chosen to end this phase of our analysis with their most recent comprehensive review on the topic, published in 2010 [hochbruck2010exponential]. This choice allows us to provide a more rigorous and detailed examination of the last 15 years of developments in the next section, which constitute a fundamental part of this work, and to which we dedicated Section 4.
@@ 3.1. The first approaches
We devote these paragraphs to the first decades of theory and implementation on ExpRK method. In particular, the first chronological reference dates back to 1958: in [hersch1958contribution], Hersch introduces the term distribution to indicate a numerical method which uses matrix exponentiation “according to the variation of constants formula”, and applies it to a semilinear ODE. Shortly thereafter, in 1960, Certaine [certaine1960solution] has started the analysis around the concept of stiffness, without explicitly using this term. The class of problems analyzed is semilinear, with a diagonal linear term, and a non-stiff nonlinear part. In the same years, Pope [pope1963exponential] proposed a manipulation technique to write a generic ODE in a semilinear form, thus introducing the easiest exponential Rosenbrock method (without using this name), and observing unconditional stability [^4] To the best of the authors’ knowledge, this is the first use of the expression “exponential method”..
The years spanning the late 1960s and early 1970s were particularly fruitful for the development of ExpRK theory. Indeed, in 1967, Lawson [lawson1967generalized] extended Pope’s results with a “Generalized RK” method. This is achieved by transforming an ODE of the form $u^{\prime}(t)=f(t,u(t))$ into one of the form $z(t)=\mathrm{e}^{-tA}u(t)$ such that $\partial f/\partial u-A$ is non-stiff (i.e. the Jacobian matrix of fAf-A has small eigenvalues), and finally applying a classical RK scheme to the transformed system. This is the basis of the integrating factor methods (IF), as opposed to exponential time differencing methods (ETD) used by Certaine and Pope [certaine1960solution, pope1963exponential], which are based on the variation of constants formula (VCF). Here, AA-stability is observed, differently from classical explicit RK methods.
In 1969, Nørsett [norsett1969stable] built a nonlinear AA-stable method by modifying the Adams-Bashforth method using VCF. The method integrates exactly $u^{\prime}(t)=-Pu(t)+T(t)$, and the author observes AA-stability. This method is of order $q+1$, where $q$ is number of past steps, and $q=0$ gives the Euler method (3), sometimes referred to as Nørsett-Euler. Concerning the analysis of stability, Guderley and Hsu [guderley1972predictor] provided in 1972 accuracy and stability analysis of the methods proposed by Certaine in 1960, considering ODEs of the form $u^{\prime}(t)=-\Lambda u(t)+f(t,u)$, with $\Lambda$ diagonal, and defining a predictor-corrector method starting from an exponential formula. They remarked that, if $\Lambda$ is not diagonal, then one needs an approximation for $\mathrm{e}^{-\Lambda t}$. Four years later, Snider [snider1976error] would correct a point overlooked by Guderley and Hsu in the truncation error analysis, although the order of accuracy does not change. In these years, Chu and Berman [chu1974exponential] also remarked that classic implicit RK method are sufficient for “not too large” ODE systems.
@@ 3.2. The rebirth
Interest in these topics seemed to fade during the 1980s, only to resurface at the end of the decade. Indeed, to the best of the authors’ knowledge, the next contribution to the theory were provided by Gallopoulos and Saad in 1989 [gallopoulos1989parallel]. The authors considered systems of the form $u^{\prime}(t)=-Au(t)+r$, that arise from discretization of linear parabolic PDEs. This translates into $A$ being typically large and sparse, and possibly stiff, and they observed that the explicit solution can be easily expressed in terms of $\mathrm{e}^{-tA}$ but that the latter is hard to compute for the typical $A$ and little changes if the solution is obtained with a time-stepping procedure. Thus, they proposed to solve the system via approximations of the exponential, either with a polynomial (Krylov method, later said to be not always good for stiff systems) or with suitable rational functions (Padé or Chebyshev approximations). [^5] The work was mostly motivated by the fact that the methods are amenable of parallelization procedures. The authors concluded by conjecturing that their methods and the possibility of parallelizing would make explicit methods popular again. We refer back to Section 2.2 for a brief recap of Krylov method.
@@
In 1996, Hochbruck et al. [hochbruck1998exponential] claimed that exponential integrators were not a novel discovery, but were considered impractical. They observed that, when AA is large, it can be harder to solve a linear system $(I-\tau A)x=v$ than to compute $\mathrm{e}^{\tau A}v$ (which is in turn as hard as computing $\varphi(\tau A)v$). This fact was proved the following by Hochbruck and Lubich [hochbruck1997krylov]. [^6] These works mark the beginning of Hochbruck’s long list of contributions on the topic. In [hochbruck1997krylov], the authors also improved Gallopoulos and Saad’s bounds on the error arising from the Krylov approximations. Here, the general class of exponential integrators was also considered with general parameters, and it was introduced as a class of methods for large systems of ODEs. The authors claimed that their result “justifies renewed interest in ODE methods that use the exponential or related functions of the Jacobian” (which in hindsight was a correct claim).
@@
In 1998, Beylkin et al. [beylkin1998new] observed that the exponential of a dense matrix can be onerous unless it is a convolution (or circular) matrix, in which case one can exploit the Fast Fourier Transform (FFT). In 2002, Cox and Matthews [cox2002exponential] proposed a more straightforward derivation of the ETD methods, compared to Beylkin’s, and provided explicit formulas for the coefficients. They derived new RK ETD methods of order 2, 3 and 4, and were the first to extend the methods to PDEs with (after discretization) nondiagonal linear part. In 2004, Koikari [koikari2005rooted] derived order conditions for ETD RK methods using the classical rooted tree analysis. To the best of the authors’ knowledge, this is the first mention of strong and weak order conditions. [^7] The role of weak order conditions will be clarified (at the latest) by Hochbruck and Ostermann in the 2005 papers as we are going to explain below. In the same year, Berland et al. [berland2005b] derived order conditions for several exponential methods (ETD, IF, CF Lie groups) for autonomous semilinear ODEs and proposed a general algorithm to build exponential integrators satisfying any given order. In 2005, Kassam and Trefethen [kassam2005fourth] introduced a new method based on contour integrals to compute the $\varphi$-functions in order to solve instability issues of ETDRK4 pointed out by Cox and Matthews [cox2002exponential]. They found that ETD methods with this modification outperform other classes of methods for PDEs, testing on a variety of models.
@@
2005 is also the year in which two fundamental contributions by Hochbruck and Ostermann were published. Firstly, in [hochbruck2005exponential] the authors developed implicit (ETD) ExpRK of collocation type, which are based on the approximation of the nonlinear term in the right-hand side with a collocation polynomial, and analyzed their order theoretically for linear and semilinear parabolic problems. They assumed that the linear operator is sectorial, requirement that holds for many typical linear operators in parabolic problems. The assumption on the nonlinear term is local Lipschitz continuity. Then, in a second paper, Hochbruck and Ostermann [hochbruck2005explicit] focused on explicit (ETD) ExpRK methods in the context of semilinear parabolic problems, dedicating additional efforts to the corresponding convergence analysis. The authors derived stiff order conditions tailored to this framework up to order $p=4$. Lastly, they proved that, in the case of constant step size, the conditions of order $p$ only have to be satisfied in a weaker form (in the sense of Koikari [koikari2005rooted]). [^8] We remark that Dujardin [dujardin2009exponential] in 2009 extended the results in Hochbruck [hochbruck2005explicit] to the Schrödinger equation (not parabolic).
@@
Still in 2005, Minchev and Wright [minchev2005review] produced a review on exponential integrators for problems for which an approximation of the Jacobian is known. Qualitatively, these are semilinear problems with non-stiff nonlinear part. The authors addressed implementation issues, that are difficulties concerning the computation of exponential and related functions, referring in particular to Moler [moler2003nineteen]. In 2006, Berland et al. [berland2005solving] comparted the performance of an IF and an ETD method of non-stiff order 4 on the Schrödinger equation. In particular, they concluded that the IF method has a generally worse performance and is more sensitive to the regularity of $V$, while the ETD method is more sensitive on the regularity of the initial condition. More importantly, they commented on the fact that the equation is not really amenable of the stiff order analysis by Hochbruck and Ostermann [hochbruck2005explicit, hochbruck2005exponential], since it is not a parabolic problem. Indeed, the tested methods have stiff orders 1 and 2 respectively, but this is not evident from the experiments.
@@
In 2008, Caliari and Ostermann [caliari2009implementation] tested interpolation on Leja points on exponential methods of Rosenbrock type (expRosenbrock) for parabolic problems. This approach seemed particularly useful since the exact Jacobian does not have a specific structure and changes at each step, and performed particularly well on problems with large advection and moderate diffusion. The authors provide order conditions for expRosenbrock methods up to order 4, remarking that they are much simpler to formulate than those for ExpRK methods. In 2009, Hochbruck et al. [hochbruck2009exponential] reformulated the expRosenbrock methods in terms of the perturbations $D^{i}_{n}:=g(U^{i}_{n})-g(u_{n})$, where $g$ is the remainder right-hand side minus the Jacobian. Since the perturbations are small vectors, applications of matrix functions to them using Krylov subspace methods require low-dimensional subspaces and therefore it would not be as computationally expensive as it was without the reformulation, where those same matrix functions would be applied to $g(u_{n})$ or $g(U^{i}_{n})$ instead. The authors provided rigorous convergence analysis up to order 4, as well as numerical tests on an advection-diffusion-reaction problem and a Schrödinger equation with time dependent potential.
@@
In 2010, Al-Mohy and Higham [al2010new] provided heuristic improvements to the original scaling and squaring algorithm behind MATLAB’s expm, motivated by the propensity for overscaling of the latter. This means that in various cases too large of a value of $s$ (number of times that one scales and squares) is needed and causes loss in floating-point arithmetic. The improvements mainly aim at addressing overscaling for upper triangular matrices, for which the problem is somehow more visible on average. The same authors, in [higham2010computing], provided a comprehensive presentation of the state of the art regarding the computation $f(A)$ where $f$ belongs to a class of functions which include the exponential. The category of methods based on similarity is represented by the Schurr-Parlett algorithm proposed in Davies [davies2003schur]. The rational approximation category is represented by Padé approximations with the improved scaling and squaring from Al-Mohy and Higham [al2010new]. The algorithms which compute $f(A)b$ without needing $f(A)$ are represented by those based on Krylov subspaces, where $f$ is only needed explicitly on a smaller Hessenberg matrix. 2010 represents a watershed moment also for this survey, as it was in that year that Hochbruck and Ostermann [hochbruck2010exponential] published the last fundamental review on exponential integrators for “two types of stiff equations”. The first type of equations are those with eigenvalues of the Jacobian having very large negative real part (the “usual definition of stiffness”) while the second is highly oscillatory problems (purely imaginary eigenvalues with large modulus).

### v3

@@ 2.3. Stability
A good starting point is provided by the paper published in 2009 by Maset and Zennaro [76], in which the authors defined the concepts of unconditional stability for ExpRK methods. Their definition is based on the concepts of contractivity and asymptotic stability for semilinear systems of the form (1) where $g$ is non-stiff. Specifically, given solutions $u_{1}$, $u_{2}$ of (1) and $\delta:=u_{1}-u_{2}$, (1) is called contractive if
@@
Maset and Zennaro [76] obtained sufficient conditions for unconditional contractivity and asymptotic stability in terms of a suitable stability function. This allows them to conclude that the ETD Euler method (3) and the two main classes of ExpRK methods for certain values of the parameters are unconditionally contractive and asymptotically stable. In 2013, the same authors [77] sharpened the aforementioned results and determined necessary and sufficient conditions for unconditional contractivity and asymptotic stability. This implies that no other ExpRK method with 2 stages, other than those already detected in [76], is unconditionally stable. They also analyzed the main methods with 3 to 5 stages, and observed that none of them is unconditionally stable either. However, since $\gamma$ is assumed to be relatively small, conditional contractivity and stability properties are also desirable.
@@
Sufficient conditions for conditional properties are provided in [77]. Without any restriction on the norm, these conditions only guarantee conditional contractivity [asymptotic stability] for those methods which are already known to be unconditionally contractive [asymptotically stable]. However, more can be said if the Euclidean norm is considered. In particular, weaker restrictions on the parameters of two-stages methods are needed in order to obtain conditional properties. The higher-order methods considered do not satisfy the sufficient conditions for general problems of the form (1), even when considering the Euclidean norm, but many satisfy conditional stability properties when restricted to certain classes of problems, such as those where AA is symmetric.
@@ 3. The historical path of ExpRK methods
In this section, we present a brief historical overview of the development and dissemination of ExpRK methods, tracing their evolution from their inception to the 2000s. We identified two major periods in which to structure our analysis. The first period begins in the late 1950s and lays the theoretical foundations that would gain traction in the following decades, up until the second half of the 1970s. The second period marks a resurgence of interest in these methods, starting in the late 1980s. Several works have already synthesized the progress made up to the early 2000s (see e.g. [79]). Therefore, we will not dwell on the details and instead refer the reader to the relevant literature. Although the 2005 works by Hochbruck and Ostermann [50, 49] represent a significant milestone in the evolution of ExpRK methods—marking the first application of these methods to semilinear parabolic problems—we have chosen to end this phase of our analysis with their most recent comprehensive review on the topic, published in 2010 [51]. This choice allows us to provide a more rigorous and detailed examination of the last 15 years of developments in the next section, which constitute a fundamental part of this work, and to which we dedicated Section 4.
@@ 3.1. The first approaches
We devote these paragraphs to the first decades of theory and implementation on ExpRK method. In particular, the first chronological reference dates back to 1958: in [41], Hersch introduces the term distribution to indicate a numerical method which uses matrix exponentiation “according to the variation of constants formula”, and applies it to a semilinear ODE. Shortly thereafter, in 1960, Certaine [17] has started the analysis around the concept of stiffness, without explicitly using this term. The class of problems analyzed is semilinear, with a diagonal linear term, and a non-stiff nonlinear part. In the same years, Pope [86] proposed a manipulation technique to write a generic ODE in a semilinear form, thus introducing the easiest exponential Rosenbrock method (without using this name), and observing unconditional stability [^4] To the best of the authors’ knowledge, this is the first use of the expression “exponential method”..
The years spanning the late 1960s and early 1970s were particularly fruitful for the development of ExpRK theory. Indeed, in 1967, Lawson [63] extended Pope’s results with a “Generalized RK” method. This is achieved by transforming an ODE of the form $u^{\prime}(t)=f(t,u(t))$ into one of the form $z(t)=\mathrm{e}^{-tA}u(t)$ such that $\partial f/\partial u-A$ is non-stiff (i.e. the Jacobian matrix of fAf-A has small eigenvalues), and finally applying a classical RK scheme to the transformed system. This is the basis of the integrating factor methods (IF), as opposed to exponential time differencing methods (ETD) used by Certaine and Pope [17, 86], which are based on the variation of constants formula (VCF). Here, AA-stability is observed, differently from classical explicit RK methods.
In 1969, Nørsett [84] built a nonlinear AA-stable method by modifying the Adams-Bashforth method using VCF. The method integrates exactly $u^{\prime}(t)=-Pu(t)+T(t)$, and the author observes AA-stability. This method is of order $q+1$, where $q$ is number of past steps, and $q=0$ gives the Euler method (3), sometimes referred to as Nørsett-Euler. Concerning the analysis of stability, Guderley and Hsu [39] provided in 1972 accuracy and stability analysis of the methods proposed by Certaine in 1960, considering ODEs of the form $u^{\prime}(t)=-\Lambda u(t)+f(t,u)$, with $\Lambda$ diagonal, and defining a predictor-corrector method starting from an exponential formula. They remarked that, if $\Lambda$ is not diagonal, then one needs an approximation for $\mathrm{e}^{-\Lambda t}$. Four years later, Snider [89] would correct a point overlooked by Guderley and Hsu in the truncation error analysis, although the order of accuracy does not change. In these years, Chu and Berman [18] also remarked that classic implicit RK methods are sufficient for “not too large” ODE systems.
@@ 3.2. The rebirth
Interest in these topics seemed to fade during the 1980s, only to resurface at the end of the decade. Indeed, to the best of the authors’ knowledge, the next contribution to the theory were provided by Gallopoulos and Saad in 1989 [35]. The authors considered systems of the form $u^{\prime}(t)=-Au(t)+r$, that arise from discretization of linear parabolic PDEs. This translates into $A$ being typically large and sparse, and possibly stiff, and they observed that the explicit solution can be easily expressed in terms of $\mathrm{e}^{-tA}$ but that the latter is hard to compute for the typical $A$ and little changes if the solution is obtained with a time-stepping procedure. Thus, they proposed to solve the system via approximations of the exponential, either with a polynomial (Krylov method, later said to be not always good for stiff systems) or with suitable rational functions (Padé or Chebyshev approximations). [^5] The work was mostly motivated by the fact that the methods are amenable of parallelization procedures. The authors concluded by conjecturing that their methods and the possibility of parallelizing would make explicit methods popular again. We refer back to Section 2.2 for a brief recap of Krylov method.
@@
In 1996, Hochbruck et al. [46] claimed that exponential integrators were not a novel discovery, but were considered impractical. They observed that, when AA is large, it can be harder to solve a linear system $(I-\tau A)x=v$ than to compute $\mathrm{e}^{\tau A}v$ (which is in turn as hard as computing $\varphi(\tau A)v$). This fact was proved the following by Hochbruck and Lubich [47]. [^6] These works mark the beginning of Hochbruck’s long list of contributions on the topic. In [47], the authors also improved Gallopoulos and Saad’s bounds on the error arising from the Krylov approximations. Here, the general class of exponential integrators was also considered with general parameters, and it was introduced as a class of methods for large systems of ODEs. The authors claimed that their result “justifies renewed interest in ODE methods that use the exponential or related functions of the Jacobian” (which in hindsight was a correct claim).
@@
In 1998, Beylkin et al. [10] observed that the exponential of a dense matrix can be onerous unless it is a convolution (or circular) matrix, in which case one can exploit the Fast Fourier Transform (FFT). In 2002, Cox and Matthews [20] proposed a more straightforward derivation of the ETD methods, compared to Beylkin’s, and provided explicit formulas for the coefficients. They derived new RK ETD methods of order 2, 3 and 4, and were the first to extend the methods to PDEs with (after discretization) nondiagonal linear part. In 2004, Koikari [58] derived order conditions for ETD RK methods using the classical rooted tree analysis. To the best of the authors’ knowledge, this is the first mention of strong and weak order conditions. [^7] The role of weak order conditions will be clarified (at the latest) by Hochbruck and Ostermann in the 2005 papers as we are going to explain below. In the same year, Berland et al. [6] derived order conditions for several exponential methods (ETD, IF, CF Lie groups) for autonomous semilinear ODEs and proposed a general algorithm to build exponential integrators satisfying any given order. In 2005, Kassam and Trefethen [57] introduced a new method based on contour integrals to compute the $\varphi$-functions in order to solve instability issues of ETDRK4 pointed out by Cox and Matthews [20]. They found that ETD methods with this modification outperform other classes of methods for PDEs, testing on a variety of models.
@@
2005 is also the year in which two fundamental contributions by Hochbruck and Ostermann were published. Firstly, in [50] the authors developed implicit (ETD) ExpRK of collocation type, which are based on the approximation of the nonlinear term in the right-hand side with a collocation polynomial, and analyzed their order theoretically for linear and semilinear parabolic problems. They assumed that the linear operator is sectorial, requirement that holds for many typical linear operators in parabolic problems. The assumption on the nonlinear term is local Lipschitz continuity. Then, in a second paper, Hochbruck and Ostermann [49] focused on explicit (ETD) ExpRK methods in the context of semilinear parabolic problems, dedicating additional efforts to the corresponding convergence analysis. The authors derived stiff order conditions tailored to this framework up to order $p=4$. Lastly, they proved that, in the case of constant step size, the conditions of order $p$ only have to be satisfied in a weaker form (in the sense of Koikari [58]). [^8] We remark that Dujardin [30] in 2009 extended the results in Hochbruck [49] to the Schrödinger equation, which is not parabolic. More recently, the same author, in collaboration with Besse and Lacroix-Violet, further developed this line of research [9] by introducing and analysing high-order exponential integrators for nonlinear Schrödinger equations, including nonautonomous problems arising from rotating Bose-Einstein condensates, with rigorous convergence and structure-preservation results.
@@
Still in 2005, Minchev and Wright [79] produced a review on exponential integrators for problems for which an approximation of the Jacobian is known. Qualitatively, these are semilinear problems with non-stiff nonlinear part. The authors addressed implementation issues, that are difficulties concerning the computation of exponential and related functions, referring in particular to Moler [82]. In 2006, Berland et al. [7] comparted the performance of an IF and an ETD method of non-stiff order 4 on the Schrödinger equation. In particular, they concluded that the IF method has a generally worse performance and is more sensitive to the regularity of $V$, while the ETD method is more sensitive on the regularity of the initial condition. More importantly, they commented on the fact that the equation is not really amenable of the stiff order analysis by Hochbruck and Ostermann [49, 50], since it is not a parabolic problem. Indeed, the tested methods have stiff orders 1 and 2 respectively, but this is not evident from the experiments.
@@
In 2008, Caliari and Ostermann [15] tested interpolation on Leja points on exponential methods of Rosenbrock type (expRosenbrock) for parabolic problems. This approach seemed particularly useful since the exact Jacobian does not have a specific structure and changes at each step, and performed particularly well on problems with large advection and moderate diffusion. The authors provide order conditions for expRosenbrock methods up to order 4, remarking that they are much simpler to formulate than those for ExpRK methods. In 2009, Hochbruck et al. [48] reformulated the expRosenbrock methods in terms of the perturbations $D^{i}_{n}:=g(U^{i}_{n})-g(u_{n})$, where $g$ is the remainder right-hand side minus the Jacobian. Since the perturbations are small vectors, applications of matrix functions to them using Krylov subspace methods require low-dimensional subspaces and therefore it would not be as computationally expensive as it was without the reformulation, where those same matrix functions would be applied to $g(u_{n})$ or $g(U^{i}_{n})$ instead. The authors provided rigorous convergence analysis up to order 4, as well as numerical tests on an advection-diffusion-reaction problem and a Schrödinger equation with time dependent potential.
@@
In 2010, Al-Mohy and Higham [1] provided heuristic improvements to the original scaling and squaring algorithm behind MATLAB’s expm, motivated by the propensity for overscaling of the latter. This means that in various cases too large of a value of $s$ (number of times that one scales and squares) is needed and causes loss in floating-point arithmetic. The improvements mainly aim at addressing overscaling for upper triangular matrices, for which the problem is somehow more visible on average. The same authors, in [42], provided a comprehensive presentation of the state of the art regarding the computation $f(A)$ where $f$ belongs to a class of functions which include the exponential. The category of methods based on similarity is represented by the Schurr-Parlett algorithm proposed in Davies [26]. The rational approximation category is represented by Padé approximations with the improved scaling and squaring from Al-Mohy and Higham [1]. The algorithms which compute $f(A)b$ without needing $f(A)$ are represented by those based on Krylov subspaces, where $f$ is only needed explicitly on a smaller Hessenberg matrix. 2010 represents a watershed moment also for this survey, as it was in that year that Hochbruck and Ostermann [51] published the last fundamental review on exponential integrators for “two types of stiff equations”. The first type of equations are those with eigenvalues of the Jacobian having very large negative real part (the “usual definition of stiffness”) while the second is highly oscillatory problems (purely imaginary eigenvalues with large modulus).